Circular No. 6916

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 K3 (LINEAR) Another comet has been found among the fast-moving asteroidal ob-

jects reported by the Lincoln Laboratory Near-Earth Asteroid Research project and included on The NEO Confirmation Page (http://cfa-

www.harvard.edu/iau/NEO/ToConfirm.html):

1998 UT  α_{2000}  δ_{2000}  m_{2}
May 23.36714  16^{h}36^{m}54.^{s}32  +2^{o}27′43′′.5  18.4

On sending the first confirmatory observations on May 24.5 UT, F. B. Zoltowski (Woomera, 0.3-m Schmidt-Cassegrain) remarked on the “poor” nature of the images, and on May 25.5 he added that they were diffuse and difficult to measure (unlike those of the six other LINEAR objects he also observed at that time). Initial orbit computations at the Minor Planet Center suggested that the orbit was retrograde. In response to our request, M. Tichý (Kleť, 0.6-m reflector), observing on May 25.9, also remarked that the object seemed slightly diffuse (∼ 8′′ coma). In better conditions on May 26.2, W. B. Offutt (Cloudcroft, 0.6-m reflector) remarked on a faint coma, ∼ 3′′ in diameter, and a faint, but discernible, tail extending ∼ 10′′ from the central condensation in p.a. ∼ 120^{o}. S. Nakano reports that A. Sugie (Dynic, 0.6-m reflector), observing on May 26.6, independently noted the object’s cometary nature, with a 6′′–8′′ coma and a 6′′ tail to the east.

The following parabolic orbital elements, computed by the undersigned from 45 observations May 23–26, are taken from MPEC 1998-K19:

T = 1998 Mar. 3.897 TT  ω = 46.^{o}986
Ω = 307.813 }2000.0
q = 3.53934 AU  i = 160.184
1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 22 16^{h}40.^{m}27 + 2^{o}15.^{′}1 2.671 3.612 154.^{o}8 6.^{o}9 17.7
27 16 27.94 + 3 00.0 2.675 3.622 155.5  6.6 17.7
June 1 16 15.67 + 3 40.3 2.693 3.631 153.9  7.1 17.8
6 16 03.66 + 4 15.5 2.726 3.642 150.3  7.9 17.8
11 15 52.10 + 4 45.0 2.771 3.653 145.4  9.1 17.8
16 15 41.13 + 5 08.9 2.829 3.664 139.9 10.3 17.9
21 15 30.87 + 5 27.2 2.897 3.676 134.0 11.5 18.0
26 15 21.40 + 5 40.3 2.976 3.689 127.9 12.6 18.0
July 1 15 12.79 + 5 48.4 3.063 3.702 121.9 13.5 18.1
6 15 05.03 + 5 52.3 3.157 3.715 115.9 14.2 18.2
1998 May 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden